OpenStudy (anonymous):

Solve -28x = 4y + 16 for y. Put your answer in the form y=mx+b with no spaces.

OpenStudy (anonymous):

Answer choices are... y=-7x-4 y=7x+4 y=-7x+4 y=7x-4

OpenStudy (eli_moses):

\[4y=-28x-16\] so \[y=-7x-4\]